the length of a rectangle is three times the width. the perimeter is 49 units. find the length.
Answer by checkley77(12844)

L=3W
P=2L+2W
49=2*3W+2W
49=6W+2W
8W=49
W=49/8
W=6.125 ANS. FOR THE WIDTH.
L=3*6.125=18.375 ANS. FOR THE LENGTH.
PROOF:
49=2*18.375+2*6.125
49=36.75+12.25
49=49